# Distributed and heterogeneous tensor–vector contraction algorithms for high performance computing

### v1

@@ Abstract
The tensor–vector contraction (TVC) is the most memory-bound operation of its class and a core component of the higher order power method (HOPM). This paper brings distributed-memory parallelization to a native TVC algorithm for dense tensors that overall remains oblivious to contraction mode, tensor splitting and tensor order. Similarly, we propose a novel distributed HOPM, namely dHOPM3, that can save up to one order of magnitude of streamed memory and is about twice as costly in terms of data movement as a distributed TVC operation (dTVC) when using task-based parallelization. The numerical experiments carried out in this work on three different architectures featuring multi-core and accelerated systems confirm that the performance of dTVC and dHOPM3 remains relatively close to the peak system memory bandwidth (50%–80%, depending on the architecture) and on par with STREAM reference values. On strong scalability scenarios, our native multi-core implementations of these two algorithms can achieve similar and sometimes even greater performance figures than those based upon state-of-the-art CUDA batched kernels. Finally, we demonstrate that both computation and communication can benefit from mixed precision arithmetic also in cases where the hardware does not support low precision data types natively.
@@
journal: Future Generation Computer Systemscopyrightcopyrightfootnotetext: ©2025. This manuscript version is made available under the CC BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/
@@ 1 Introduction
Tensors can be considered as multidimensional arrays that store data in a certain manner according to multiple (i.e., multilinear) attributes. Exploiting them is of great importance since it allows to extract patterns inherently present in such datasets. It is the field of multilinear algebra that defines tensor operations and algorithms such as products, transformations, and decompositions to name but a few [1]. A major kernel is the tensor contraction, which includes the following operations: (i) tensor–tensor contraction (TTC), (ii) tensor–matrix contraction (TMC), and (iii) tensor–vector contraction (TVC), being all of them core components in widely used tensor algorithms: the Khatri-Rao product [2], the higher order orthogonal iteration (HOOI) algorithm [3] that computes the truncated Tucker decomposition [4], the alternating least squares (ALS) algorithm for the canonical polyadic decomposition (CPD) [5], or the higher order power method (HOPM) [3].
@@
The actual computation inside a TVC kernel can be achieved in different ways, e.g. via the “looped” and “unfolded” algorithms. Both of them appeal to the use of MVM operations, which requires the conversion of the input tensor into its matricized form before the actual contraction takes place. While the looped variant recognizes the input tensor as a series of contiguous matrices and carries out several MVMs, the unfolded counterpart first needs to reorganize the tensor in memory, at the expense of additional computation and data movement, in order to assemble a single large matrix and conduct one unique MVM. The popularity of these two approaches resides in the fact that they rely upon a broadly available and optimized BLAS level 2 routine: dgemv(). Nevertheless, TVC is far more susceptible to mode-awareness —thus indicating that performance is strongly influenced by the contraction mode $k$— than a regular second-order contraction: either vector–matrix ($k=0$) or matrix–vector ($k=1$) multiplication. Indeed, in a shared-memory parallel context, it may not be possible to evenly distribute MVMs among available threads in the looped TVC algorithm, resulting in suboptimal performance [6]. For this reason, authors have recently begun paying attention to mode-oblivious TVC procedures [6, 7, 8].
@@
    Developing distributed-memory, highly optimized TVC and HOPM algorithms and rendering them publicly available in an open-source library.
@@
The remainder of this paper is organized as follows. Sections 2 and 3 are dedicated to describing basic aspects about tensors and giving some related work. In Section 4 we detail our distributed TVC and HOPM algorithms. We carry out our numerical experiments on Section 5 to assess the performance of the aforementioned algorithms and, finally, Section 6 presents the final conclusions and future work.
@@ 2 Background
We briefly introduce some basic concepts about tensors taken mainly from Kolda & Bader [1] and reuse the notation introduced in our previous work [8]. The font shapes $x$, $\mathbf{x}$, $\mathbf{X}$ and $\mathcal{X}$ are used to refer to scalars, vectors, matrices, and tensors, respectively. We interpret dense tensors as multidimensional arrays stored in system memory following a last-order, nonhierarchical storage layout (see Fig. 1) which is indeed equivalent to the multidimensional array ordering used by the C/C++ programming languages. Similarly to these languages, we adopt a zero-based indexing.
@@
For the actual computation of the previous TVC it is practical to reinterpret the input tensor $\mathcal{A}$ in matrix form. By defining $u=\prod_{i=0}^{k-1}n_{i}$ and $v=\prod_{i=k+1}^{d-1}n_{i}$ one can form the matricized form of such tensor $\mathbf{A}^{uv\times n_{k}}$ and its transpose $\mathbf{A}^{n_{k}\times uv}$. Now a looped TVC algorithm can be built upon a series of classical matrix–vector multiplications over contiguous subsets of $\mathbf{A}$. The kind of operation depends on the memory layout and, for the last-order arrangement adopted herein, one single matrix–vector multiplication, $\mathcal{Y}=\mathbf{A}^{u\times n_{k}}\times_{k}\mathbf{x}$ is required for the last contraction mode $k=d-1$; otherwise, $u$ independent, left-hand sided vector–matrix multiplications, $\mathcal{Y}=\bigsqcup_{u}\mathbf{x}^{\intercal}\times_{k}\mathbf{A}^{n_{k}\times v}$, are necessary for the remaining modes $k<d-1$. Note that $u=1$ for $k=0$ and $v=1$ for $k=d-1$.
@@
The looped algorithm described above has been widely adopted by tensor libraries and can be parallelized at two levels: (i) the matrix–vector (or vector–matrix) multiplication itself and/or (ii) the $u$ independent multiplications for $10<k<d-$ 1. HPC libraries such as Intel oneAPI MKL and NVIDIA cuBLAS already provide optimized (i.e., batched) functions [10] for subsequent MVMs: cblas_gemv_batch_ strided() and cublas_gemvStridedBatched(), respectively. Nevertheless, all these efforts do not prevent this looped algorithm from being exposed to mode-aware, suboptimal performance and, therefore, it becomes necessary to find alternative approaches to the looped TVC.
3 Related work
There are numerous examples of tensor–vector contraction algorithms, the majority of them being based upon BLAS level 2 and 3 routines [11, 12]. Consequently, it is not rare that such algorithms heavily exploit HPC libraries such as Intel oneMKL, OpenBLAS [13], as well as BLIS [14] and LIBXSMM [15], which provide better performance over repetitive small matrix–matrix multiplications. One of the main challenges of TVC, besides of its memory-bound nature, is its performance reliance on the contraction mode. As a result, its global efficiency can be compromised even with highly optimized BLAS routines. Pawłowski et al. introduced a sequential, mode-oblivious native TVC algorithm based on a Morton-ordered memory layout [6] and later parallelized it on shared-memory systems [7]. In this respect, we recently proposed a shared-memory, mode-oblivious native TVC algorithm for nonhierarchically stored dense tensors [8] that naturally distributes the column space of $\mathbf{A}^{n_{k}\times uv}$ among CPU cores thereby resulting in a specialized, nonstandard BLAS level 2 routine.
With the advent of big data analysis, it has become necessary to perform tensor calculations on large scale, distributed systems. Existing tensor libraries typically fall within two main categories. In the first category, the actual distribution takes place at the filesystem level via the “mapReduce” programming paradigm popularized by Hadoop, e.g., GigaTensor [16] and BIGtensor [17], as well as Spark, e.g. CSTF [18]. In general, these distributed implementations do not support higher order tensors and require additional tensor unfolding operations owing to the explicit parallel distribution which, in the end, result in filesystem overhead. The second category is more interesting from a HPC perspective with applications making extensive use of the MPI library [19]. The Algebraic Programming (ALP) framework [^1] Visit: https://algebraic-programming.github.io/ leverages explicit algebraic annotations to computations for optimization and auto-parallelization, both shared- and distributed-memory. ALP/GraphBLAS and the wider GraphBLAS efforts promote the paradigm for generalised sparse linear algebra [20, 21], whereas ALP/Dense adds dense linear algebra support [22]. Two remarkable tensor frameworks for distributed systems are CTF [23] and tiledArray [24], both achieving similar levels of parallel performance. CTF aims at minimizing communication cost, rather than necessarily achieving best performance. Its flexibility for multidimensional tensor splitting implies data redistribution (copying and transposition), hence giving rise to very low flop-to-byte ratio algorithms that may not achieve memory bandwidth peak performance. Ultimately, these HPC tensor libraries resort to batched BLAS primitives (e.g. from LAPACK [25], ScaLAPACK [26] and others cited above) with the aforementioned implications on looped TVC performance. All things considered, the main advantage offered by the aforementioned frameworks resides in the simplicity and generality of expressing tensor operations via a domain-specific language (DSL).
To the best of our knowledge, there are no previous works focusing on the performance of a mode-oblivious TVC algorithm on distributed systems using higher order dense tensors and evaluating the effects of tensor splitting on bandwidth from both analytical and numerical perspectives. Moreover, no other BLAS level 2 TVC algorithm with oblivious properties has been proposed for hybrid computing, or compared against state-of-the-art tensor libraries for distributed systems, or tailored GPU kernels for heterogeneous computing. In this regard, the present work showcases the importance of seeking near theoretical bandwidth performance during distributed TVC operations, making special emphasis on a real-world tensor operation: the distributed higher order power method. We hope that our contribution will be used as an example of future development around HPC libraries oriented towards tensor operations on distributed and heterogeneous systems.
4 Algorithm description
@@
This section firstly discusses the particularities of tensor splitting on distributed systems during tensor–vector contractions and, secondly, it proposes an optimized version of the distributed higher order power method.
@@ 4.1 Distributed-memory tensor–vector contraction
To allow for distributed-memory TVC computations, one can split the given input tensor $\mathcal{A}$ along one or more dimensions. Such splitting also influences the output tensor $\mathcal{Y}$ that is distributed in a similar manner. The input vector $\mathbf{x}$ that participates in the contraction can be harmlessly duplicated among the $p$ distributed processes since, in general, we can assume that $uv\gg n_{k}$. In this work, we only consider one-dimensional cuts of both the input and output tensors for three main reasons: (i) it yields minimum communication [7], (ii) it gives best computational performance as it does not incur additional tensor unfolding operations, and (iii) it greatly simplifies tensor distribution and reassembly. The major disadvantage of 1D splitting resides in the fact that the maximum number of processes that can be used is determined by $\max(n_{i})$. Such processes can be bound to NUMA nodes, GPU devices or even entire compute nodes thereby maximizing the potential for parallelism.
Figure 1: One-dimensional optimal splittings ($s=0,1,2$) using up to three processes ($p=3$) and a third-order tensor $\mathcal{A}^{2\times 3\times 4}$ stored in last-order, nonhierarchical memory layout.
[…]
We now proceed with the tensor–vector contraction algorithm in terms of both intra- and inter-process data movement. A general expression for the TVC operation can be written as
@@
when the contraction mode and the splitting dimension coincide ($k=s$). This second case remains suboptimal because each distributed TVC operation yields $p$ subtensors with the exact same size as the global result $N/n_{k}$, which translates into more local computations and hence streamed memory compared to the first case. Additionally, Eq. (2) demands a collective summation of all these equally-sized subtensors, element by element, to get the global tensor, hence the sum symbol. In contrast, Eq. (1) incurs much less communication because each subtensor only consists of $N[n_{s}/p]/(n_{k}n_{s})\approx N/(n_{k}p)$ elements. Since $k=s$ leads to further communication as well as computation, one should avoid contracting tensors along their splitting dimension as much as possible. What is more important, it is strongly encouraged to work with the distributed output tensors and only construct the global tensor when strictly needed because communication is slow and available memory may also be an issue. Finally, scalars $\alpha$ and $\beta$ are typical parameters used in this type of BLAS-like functions and $\mathbf{x}^{(p)}$ in Eq. (2) indicates that the contraction operates on a disjoint subset of $\mathbf{x}$ instead of the entire vector. Getting back to the previous matrix analogy, Eq. (2) is identical to a series of vector–matrix multiplications applied to subsets of both $\mathbf{x}$ and $\mathbf{A}$, where the matrix is distributed along rows.
@@
An important aspect related to the global construction of the output tensor is the contiguity of the elements being assembled. As already mentioned, using a last-order layout with $s=0$ simply requires to contiguously gather all rows from each process in order to build the global tensor. On the other hand, for $s>0$ it becomes necessary to communicate fragments of data from each distributed tensor and interleave them inside the global tensor. To this end, one can resort to the matricized view of $\mathcal{Y}^{(p)}$ and $\mathcal{Y}$, both split along the $s-1$ dimension, to determine the groups of columns to be collected from each process. This results in $w=\prod_{i=0}^{s-2}n_{i}$ communication messages per process compared to a single one for $s=0$ (likewise, $w=1$ for $s=1$) for gathering a total of $N/n_{k}$ elements. Therefore, increasing values of $s$ are expected to negatively impact parallel performance. However, this adversity can be overcome by trading messages for further data movement: firstly, local memory requirements are duplicated to hold two global tensors; secondly, one collective per process is employed to gather all the distributed tensors consecutively in memory inside the first global (disjoint) buffer; and, lastly, columns are copied in groups to the second buffer in order to compose the global (joint) tensor thereby incurring in the additional movement of $N/n_{k}$ elements within local memory.
4.2 Distributed-memory higher order power method
The higher order power method is a generalization of the well-known power iteration algorithm applied to matrices and is employed to find the best rank-1 approximation of a tensor [3]. Given a $d$-order input tensor and a set of $d$ compatible vectors, the HOPM performs $d$ external iterations (one per dimension) within which $d-1$ TVC operations are carried out consecutively (omitting precisely the contraction along that specific mode).

### v2

@@ Abstract
The tensor–vector contraction (TVC) is the most memory-bound operation of its class and a core component of the higher-order power method (HOPM). This paper brings distributed-memory parallelization to a native TVC algorithm for dense tensors that overall remains oblivious to contraction mode, tensor splitting and tensor order. Similarly, we propose a novel distributed HOPM, namely dHOPM3, that can save up to one order of magnitude of streamed memory and is about twice as costly in terms of data movement as a distributed TVC operation (dTVC) when using task-based parallelization. The numerical experiments carried out in this work on three different architectures featuring multi-core and accelerators confirm that the performances of dTVC and dHOPM3 remain relatively close to the peak system memory bandwidth (50%–80%, depending on the architecture) and on par with STREAM benchmark figures. On strong scalability scenarios, our native multi-core implementations of these two algorithms can achieve similar and sometimes even greater performance figures than those based upon state-of-the-art CUDA batched kernels. Finally, we demonstrate that both computation and communication can benefit from mixed precision arithmetic also in cases where the hardware does not support low precision data types natively.
@@
journal: Future Generation Computer Systems
@@ 1 Introduction
Tensors can be considered as multidimensional arrays that store data in a certain manner according to multiple (i.e., multilinear) attributes. Exploiting them is of great importance since it allows to extract patterns inherently present in such datasets. It is the field of multilinear algebra that defines tensor operations and algorithms such as products, transformations, and decompositions to name but a few [1]. A major kernel is the tensor contraction, which includes the following operations: (i) tensor–tensor contraction (TTC), (ii) tensor–matrix contraction (TMC), and (iii) tensor–vector contraction (TVC), being all of them core components in widely used tensor algorithms: the Khatri-Rao product [2], the higher-order orthogonal iteration (HOOI) algorithm [3] that computes the truncated Tucker decomposition [4], the alternating least squares (ALS) algorithm for the canonical polyadic decomposition (CPD) [5], or the higher-order power method (HOPM) [3].
@@
The actual computation inside a TVC kernel can be achieved in different ways, e.g. via the “looped” and “unfolded” algorithms. Both of them appeal to the use of MVM operations, which requires the conversion of the input tensor into its matricized form before the actual contraction takes place. While the looped variant recognizes the input tensor as a series of contiguous matrices and carries out several MVMs, the unfolded counterpart first needs to reorganize the tensor in memory, at the expense of additional data movement, in order to assemble a single large matrix and conduct one unique MVM. The popularity of these two approaches resides in the fact that they rely upon a broadly available and optimized BLAS level 2 routine: gemv. Nevertheless, TVC is far more susceptible to mode-awareness —thus indicating that performance is strongly influenced by the contraction mode $k$— than a regular second-order contraction: either vector–matrix ($k=0$) or matrix–vector ($k=1$) multiplication. Indeed, in a shared-memory parallel context, it may not be possible to evenly distribute MVMs among available threads in the looped TVC algorithm, resulting in suboptimal performance [6]: up to 18% lower performance and up to 71% higher variability across contraction modes. For this reason, authors have recently begun paying attention to mode-oblivious TVC procedures [6, 7, 8].
@@
    Developing distributed-memory, highly optimized TVC and HOPM algorithms and rendering them publicly available in an open-source library [10].
@@
The remainder of this paper is organized as follows. Sections 2 and 3 are dedicated to giving some related work and describing basic aspects about tensors. In Section 4 we detail our distributed TVC and HOPM algorithms. We carry out our numerical experiments in Section 5 to assess the performance of the aforementioned algorithms and, finally, Section 6 presents the final conclusions and future work.
2 Related work
There are numerous examples of tensor–vector contraction algorithms, the majority of them being based upon BLAS level 2 and 3 routines [11, 12]. Consequently, it is not rare that such algorithms heavily exploit HPC libraries such as Intel oneMKL, OpenBLAS [13], BLIS [14], as well as LIBXSMM [15] which provides better performance over repetitive small matrix–matrix multiplications. One of the main challenges of TVC, besides of its memory-bound nature, is its performance reliance on the contraction mode. As a result, its global efficiency can be compromised even with highly optimized BLAS routines. Pawłowski et al. introduced a sequential, mode-oblivious TVC algorithm based on the Morton-ordered memory layout [6] and later parallelized it on shared-memory systems [7]. In this respect, we recently proposed a shared-memory, mode-oblivious native TVC algorithm for nonhierarchically stored dense tensors [8] that naturally distributes the column space of the matricized view of a tensor among CPU cores thereby resulting in a specialized, nonstandard BLAS level 2 routine.
With the advent of big data analysis, it has become necessary to perform tensor calculations on large scale, distributed systems. Existing tensor libraries typically fall within two main categories. In the first category, the actual distribution takes place at the filesystem level via the “mapReduce” programming paradigm popularized by Hadoop, e.g. GigaTensor [16] and BIGtensor [17], as well as Spark, e.g. CSTF [18]. In general, these distributed implementations do not support higher-order tensors and require additional tensor unfolding operations owing to the explicit parallel distribution which, in the end, result in filesystem overhead. The second category is more interesting from a HPC perspective with applications making extensive use of the MPI library [19]. The Algebraic Programming (ALP) framework [^3] Visit: https://algebraic-programming.github.io/ leverages explicit algebraic annotations to computations for optimization and auto-parallelization, both shared- and distributed-memory. ALP/GraphBLAS and the wider GraphBLAS efforts promote the paradigm for generalized sparse linear algebra [20, 21], whereas ALP/Dense adds dense linear algebra support [22]. Remarkable examples of tensor frameworks for distributed systems are CTF [23] and tiledArray [24], both achieving similar levels of parallel performance, as well as Deinsum/DaCe [25]. CTF aims at minimizing communication cost, rather than necessarily achieving best performance. Its flexibility for multidimensional tensor splitting implies data redistribution (copying and transposition), hence giving rise to very low flop-to-byte ratio algorithms that may not achieve memory bandwidth peak performance. Both CTF and tiledArray resort to batched BLAS primitives (e.g. from LAPACK [26], ScaLAPACK [27] and others cited above) with the aforementioned implications on looped TVC performance. All things considered, the main advantage offered by these two frameworks resides in the simplicity and generality of expressing tensor operations via a domain-specific language (DSL). On the other hand, Deinsum translates Python code into high performance binaries, derives data movement-optimal tiling, generates corresponding distributed schedules, and can optimize the performance of local computations by increasing their arithmetic intensity, demonstrating significant speedups over CTF.
To the best of our knowledge, there are no previous works focusing on the performance of a mode-oblivious TVC algorithm on distributed systems using higher-order dense tensors and evaluating the effects of tensor splitting on bandwidth from both analytical and numerical perspectives. Moreover, no other BLAS level 2 TVC algorithm with oblivious properties has been proposed for hybrid computing, or compared against state-of-the-art tensor libraries for distributed systems, or tailored GPU kernels for heterogeneous computing. In this regard, the present work showcases the importance of seeking near theoretical bandwidth performance during distributed TVC operations, making special emphasis on a real-world tensor operation: the distributed higher-order power method. We hope that our contribution will be used as an example of future development around HPC libraries oriented towards tensor operations on distributed and heterogeneous systems.
@@ 3 Background
We briefly introduce some basic concepts about tensors taken mainly from Kolda & Bader [1] and reuse the notation introduced in our previous work [8]. The font shapes $x$, $\mathbf{x}$, $\mathbf{X}$ and $\mathcal{X}$ refer to scalars, vectors, matrices, and tensors, respectively. We interpret dense tensors as multidimensional arrays stored in system memory following a last-order, nonhierarchical storage layout (see Fig. 1) which is indeed equivalent to the multidimensional array ordering used by the C/C++ programming languages. Similarly to these languages, we adopt a zero-based indexing.
@@
For the actual computation of the previous TVC it is practical to reinterpret the input tensor $\mathcal{A}$ in matrix form. By defining $u=\prod_{i=0}^{k-1}n_{i}$ and $v=\prod_{i=k+1}^{d-1}n_{i}$ one can build the matricized form of such tensor $\mathbf{A}^{uv\times n_{k}}$ and its transpose $\mathbf{A}^{n_{k}\times uv}$. Now a looped TVC algorithm can be built upon a series of classical matrix–vector multiplications over contiguous subsets of $\mathbf{A}$. The kind of operation depends on the memory layout and, for the last-order arrangement adopted herein, one single matrix–vector multiplication, $\mathcal{Y}=\mathbf{A}^{u\times n_{k}}\times_{k}\mathbf{x}$ is required for the last contraction mode $k=d-1$; otherwise, $u$ independent, equally-sized, left-hand sided vector–matrix multiplications of the form $\mathcal{Y}=\mathbf{x}^{\intercal}\times_{k}\mathbf{A}^{n_{k}\times v}$ are necessary for the remaining modes $k<d-1$. Note that $u=1$ for $k=0$ and $v=1$ for $k=d-1$.
@@
The sequential looped algorithm described above has been widely adopted by tensor libraries and can be parallelized at two levels: (i) the matrix–vector (or vector–matrix) multiplication itself and/or (ii) the $u$ independent multiplications for $10<k<d-$ 1. HPC libraries such as Intel oneMKL and NVIDIA cuBLAS already provide optimized (i.e., batched) functions [28] for subsequent MVMs: cblas_gemv_batch_strided for the former and cublas_gemvStridedBatched for the later. Nevertheless, all these efforts do not prevent this looped algorithm, either in its sequential or parallel version, from being exposed to mode-aware, suboptimal performance and, therefore, it becomes necessary to find alternative approaches to the looped TVC.
The higher-order power method is a generalization of the well-known power iteration algorithm applied to matrices and is employed to find the best rank-1 approximation of a tensor [3]. Given a $d$-order input tensor $\mathcal{A}$ and a set of $d$ compatible vectors $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\ldots$, $\mathbf{x}_{d-1}$, the HOPM performs $d$ external iterations within which $d-1$ TVC operations are carried out consecutively omitting, precisely, the contraction along the external mode. This results in a total of $d(d-1)$ tensor contractions, which renders the HOPM algorithm an excellent benchmark for testing the performance of different TVC implementations. We refer the reader to Refs. [3, 6] for a canonical representation of the HOPM algorithm.
It is worth mentioning that the tensor operations described in this section refer basically to sequential algorithms. Shared-memory, parallel implementations of TVC are discussed by Pawłowski et al. [7] and can also be found in our previous work [8]. Distributed-memory implementations are addressed in the next section.
4 Distributed algorithms
@@
This section firstly introduces the dTVC and discusses the particularities of tensor splitting on distributed systems during tensor–vector contractions and, secondly, it proposes an optimized version of the dHOPM. Finally, it comments on the data-flow parallelization strategy adopted for the dHOPM based on task annotations.
@@ 4.1 Distributed-memory tensor–vector contraction
To allow for distributed-memory TVC computations, one can split the given input tensor $\mathcal{A}$ along one or more dimensions. Such splitting also influences the output tensor $\mathcal{Y}$ that is distributed in a similar manner. The input vector $\mathbf{x}$ that participates in the contraction can be harmlessly duplicated among the $p$ distributed processes since, in general, we can assume that $uv\gg n_{k}$. In this work, we only consider one-dimensional cuts of both the input and output tensors for three main reasons: (i) it yields minimum communication [7], (ii) it gives best computational performance as it does not incur additional tensor unfolding operations, and (iii) it greatly simplifies tensor distribution and reassembly. The major disadvantage of 1D splitting resides in the fact that the maximum number of processes that can be used is determined by $\max(n_{i})$. Such processes can be bound to CPU cores, NUMA nodes, GPU devices or even entire compute nodes thereby maximizing the potential for parallelism.
Figure 1: One-dimensional optimal splittings ($s=0,1,2$) using up to three processes ($p=3$) for a third-order tensor $\mathcal{A}^{2\times 3\times 4}$ stored in last-order, nonhierarchical memory layout.
Let us define the contraction mode $k$ and splitting dimension $s$. Since we store tensors in last-order layout, a splitting along the first dimension of $\mathcal{A}$ guarantees the contiguity between elements of the resulting $\mathcal{A}^{(p)}$ distributed subtensors: $\mathcal{A}=\bigsqcup_{p}\mathcal{A}^{(p)}$ with $\mathcal{A}^{(p)}\in\mathbb{R}^{[n_{0}/p]\times n_{1}\times\ldots\times n_{d-1}}$. In other words, the distributed tensors can be reassembled into a global tensor straightforwardly as shown in Fig. 1 with $s=0$. The symbol $\bigsqcup_{p}$ indicates the disjoint union of $p$ tensors and the brackets in $[n_{0}/p]$ refer to the optimal division between $n_{0}$ and $p$. Usually, it corresponds to the ceiling division $\lceil n_{0}/p\rceil$ although, for vectorization purposes, we employ a heuristic technique that seeks to promote quotients that are multiple of the vector length (e.g., 512 bits). This may override (i.e., lower) the value of $p$ for an optimal splitting; for instance, Fig. 1 with $s=2$ results in $[4/3]\rightarrow 4/2$, that is $p\rightarrow 2$, so that only two of the three requested processes are actually employed. Following the same reasoning, a cut along the last dimension, $[n_{d-1}/p]$, results in the most indirect tensor distribution and later assembly: in the particular case of a second-order tensor, this corresponds to splitting along columns ($s=1$) a matrix stored in row-major memory layout.
We now proceed with the tensor–vector contraction algorithm in terms of both intra- and inter-process data movement. A general, BLAS-like expression for the distributed TVC operation with typical scalars $\alpha$ and $\beta$ can be written for $k\neq s$ as
@@
when the contraction mode and the splitting dimension coincide ($k=s$). This second case remains suboptimal because each distributed TVC operation yields $p$ subtensors with the exact same size $N/n_{k}$ as the global result, which translates into more local computations and hence streamed memory compared to the first case. Additionally, Eq. (2) demands a collective summation of all these equally-sized subtensors, element by element, to get the global tensor, hence the sum symbol. In contrast, Eq. (1) incurs much less communication because each subtensor only consists of approximately $N[n_{s}/p]/(n_{k}n_{s})\approx N/(n_{k}p)$ elements. Since $k=s$ leads to further communication as well as computation, one should avoid contracting tensors along their splitting dimension as much as possible. What is more important, it is strongly encouraged to work with the distributed output tensors and only construct the global tensor when strictly needed because communication is slow and available memory may also be an issue. Finally, $\mathbf{x}^{(p)}$ in Eq. (2) indicates that the contraction operates on a disjoint subset of $\mathbf{x}$ instead of the entire vector. Getting back to the previous matrix analogy, Eq. (2) is identical to a series of matrix–vector multiplications applied to subsets of both $\mathbf{x}$ and $\mathbf{A}$, where the matrix is distributed along columns ($k=s=1$) or a series of vector–matrix multiplications when the matrix is distributed along rows ($k=s=0$).
@@
An important aspect related to the global construction of the output tensor is the contiguity of the elements being assembled. As previously shown in Fig. 1, using the last-order layout with $s=0$ simply requires to contiguously gather all rows from each process in order to build the global tensor. On the other hand, for $s>0$ it becomes necessary to communicate fragments of data from each distributed tensor and interleave them inside the global tensor. To this end, one can resort to the matricized views of $\mathcal{Y}^{(p)}$ and $\mathcal{Y}$, both split along the $s-1$ dimension, to determine the groups of columns to be collected from each process. This results in $w=\prod_{i=0}^{s-2}n_{i}$ communication messages per process compared to a single one for $s=0$ (likewise, $w=1$ for $s=1$) for gathering a total of $N/n_{k}$ elements. Therefore, increasing values of $s$ are expected to negatively impact parallel performance. However, this adversity can be overcome by trading messages for further data movement: firstly, local memory requirements are duplicated to hold two global tensors; secondly, one collective per process is employed to gather all the distributed tensors consecutively in memory inside the first global (disjoint) buffer; and, lastly, columns are copied in groups to the second buffer in order to compose the global (joint) tensor thereby incurring in the additional movement of $N/n_{k}$ elements within local memory.
4.2 Distributed-memory higher-order power method